# Location-Agnostic Channel Knowledge Map Construction for Dynamic Scenes

### 10 Mar 2026 v1

@@ III-A Construction Pipeline
where $\mathbf{\hat{H}}^{\mathrm{D}}_{i}(f,t)$ and $\mathbf{H}^{\mathrm{D}}_{i}(f,t)$ represent the predicted and ground truth CSI of subcarrier frequency $f$ within coherence block $t$ in the $i$-th sample, respectively, with $N$ and $N_{b}$ denoting the total number of data samples and coherence blocks, respectively. In particular, this pipeline comprises three major stages:
@@ IV Simulation Results
To evaluate the proposed methods, a novel synthetic channel dataset is created in outdoor dynamic scenes via ray-tracing. As shown in Fig. 5, we consider a real-world campus environment whose 3D geometry is downloaded from OpenStreetMap [^2] https://www.openstreetmap.org/. The scene is imported into Blender for material parameterization and then fed into the Sionna ray-tracing platform [3] to generate ground-truth MIMO-OFDM channels. To simulate environment dynamics, we randomly reposition dynamic scatterers in every coherence block, where pedestrians are modeled as dry dielectric boxes ($0.5\,\text{m}\times 0.5\,\text{m}\times 1.8\,\text{m}$) and vehicles as metal boxes ($2\,\text{m}\times 4\,\text{m}\times 1.6\,\text{m}$). The OFDM bandwidth is $20$ MHz, split into $64$ sub-carriers with $312.5$ kHz spacing, where $52$ sub-carriers are used for channel measurement. The central frequencies are set to $6.715$ GHz (uplink) and $6.765$ GHz (downlink). Both BS and UE employ uniform planar arrays (UPAs) with discrete Fourier transform (DFT) codebooks. We consider BS antenna counts $N_{t}\in\{8,16,32,64,128\}$ and UE antenna counts $N_{r}\in\{1,4\}$. For each MIMO configuration $(N_{t},N_{r})$, $20$ coherence blocks are generated, with $2,500$ UEs randomly dropped per block.
We implement our LAD-CKM and all baselines using the PyTorch platform, with a $2.60$ GHz Intel(R) Xeon(R) Platinum 8350C CPU, and an NVIDIA GeForce RTX 4090 GPU. To train LAD-CKM, the Adam optimizer with an initial learning rate of $5\times 10^{-5}$ and a batch size of $2$ is employed. The learning rate is adjusted over time using the ReduceLROnPlateau scheduler with a patience of $10$ and a reduction factor of $0.9$. We compare the effective data rate performance of the following schemes:
@@
  • LAD-CKM ($\rho=0.125$): The proposed scheme with $12.5\%$ pilot overhead, where pilots are transmitted on every-$4$th subcarrier and every-22nd antenna.

### 17 Mar 2026 v2

@@ III-A Construction Pipeline
where $\mathbf{\hat{H}}^{\mathrm{D}}_{i}(f,t)$ and $\mathbf{H}^{\mathrm{D}}_{i}(f,t)$ represent the predicted and ground truth CSI of subcarrier frequency $f$ within coherence block $t$ in the $i$-th sample, respectively, with NN and $N_{b}$ denoting the total number of data samples and coherence blocks, respectively. In particular, this pipeline comprises three major stages:
@@ IV Simulation Results
To evaluate the proposed methods, a novel synthetic channel dataset is created in outdoor dynamic scenes via ray-tracing. As shown in Fig. 5, we consider a real-world campus environment whose 3D geometry is downloaded from OpenStreetMap [^2] https://www.openstreetmap.org/. The scene is imported into Blender for material parameterization and then fed into the Sionna ray-tracing platform [3] to generate ground-truth MIMO-OFDM channels. To simulate environment dynamics, we randomly reposition dynamic scatterers in every coherence block, where pedestrians are modeled as dry dielectric boxes ($0.5\,\text{m}\times 0.5\,\text{m}\times 1.8\,\text{m}$) and vehicles as metal boxes ($2\,\text{m}\times 4\,\text{m}\times 1.6\,\text{m}$). The OFDM bandwidth is 2020 MHz, split into $64$ sub-carriers with $312.5$ kHz spacing, where $52$ sub-carriers are used for channel measurement. The central frequencies are set to $6.715$ GHz (uplink) and $6.765$ GHz (downlink). Both BS and UE employ uniform planar arrays (UPAs) with discrete Fourier transform (DFT) codebooks. We consider BS antenna counts $N_{t}\in\{8,16,32,64,128\}$ and UE antenna counts $N_{r}\in\{1,4\}$. For each MIMO configuration $(N_{t},N_{r})$, 2020 coherence blocks are generated, with $2,500$ UEs randomly dropped per block.
We implement our LAD-CKM and all baselines using the PyTorch platform, with a $2.60$ GHz Intel(R) Xeon(R) Platinum 8350C CPU, and an NVIDIA GeForce RTX 4090 GPU. To train LAD-CKM, the Adam optimizer with an initial learning rate of $5\times 10^{-5}$ and a batch size of 22 is employed. The learning rate is adjusted over time using the ReduceLROnPlateau scheduler with a patience of $10$ and a reduction factor of $0.9$. We compare the effective data rate performance of the following schemes:
@@
  • LAD-CKM ($\rho=0.125$): The proposed scheme with $12.5\%$ pilot overhead, where pilots are transmitted on every-44th subcarrier and every-22nd antenna.